A boy forgot his PIN CODE which is composed of 5 digits, but luckily he remembered some hints how to remember the password. Here are those clues.
1. First digit is equal to the square of the second digit.
2. Second plus the third is equal to 10.
3. Fourth digit is equal to second plus 1.
4. Fifth plus third digit make 14; and
5. Sum of all the five digits of the PIN is 30.
What now is the pin code?

Answer by stanbon(75887)

1. First digit is equal to the square of the second digit.
2. Second plus the third is equal to 10.
3. Fourth digit is equal to second plus 1.
4. Fifth plus third digit make 14; and
5. Sum of all the five digits of the PIN is 30.
What now is the pin code?
---------------------------------------
Equations:
f = s^2
s + d = 10
t = s+1
h + d = 14
f + s + d + t + h = 30
----
s^2 + s + (10-s) + (s+1) + (14-(10-s)) = 30
---
s^2 +2s +15 = 30
s^2 + 2s - 15 = 0
Factor:
(s-5)(s+3) = 0
Positive solution:
s = 5
-----
first = 25 (note: that is not a single digit)
second = 5
third = 10-s = 5
fourth = 6
fifth = 9
===============
Cheers,
Stan H.